### v2

# EtCon: Edit then Consolidate for Reliable Knowledge Editing

Abstract
Knowledge editing aims to update specific facts in large language models (LLMs) without full retraining. Prior efforts sought to tune the knowledge layers of LLMs, achieving improved performance in controlled, teacher-forced evaluations. However, they still encounter challenges in real-world autoregressive generation scenarios, which greatly limit their practical applicability. Our empirical analysis reveals two issues: (1) Most methods degrade pre-trained capabilities after injecting new knowledge; (2) They may exhibit a discrepancy between stored parametric knowledge and inference-time autoregressive generation behavior. To this end, we propose EtCon, an edit-then-consolidate paradigm that couples targeted edits with post-edit consolidation. Specifically, our framework comprises two stages: (1) Targeted Proximal Supervised Fine-Tuning (TPSFT) performs a constrained targeted edit to update parametric knowledge while controlling policy drift. (2) Group Relative Policy Optimization (GRPO) consolidates the edit by aligning autoregressive trajectories with the intended fact. Extensive experiments demonstrate that our EtCon improves editing reliability and real-world generalization, while better preserving pre-trained capabilities.
[…]
1 Introduction
Large language models (LLMs) have demonstrated unprecedented capabilities across numerous tasks (Wang et al., 2025; Yang et al., 2025a; He et al., 2025; Liu et al., 2025b; Rosen et al., 2025; Shmatko et al., 2025). However, as the world continuously evolves, static pre-trained knowledge quickly becomes outdated (Zheng et al., 2025a). While continual pre-training can refresh knowledge, it is computationally expensive and impractical to perform frequently at scale (Mitchell et al., 2022), motivating knowledge editing methods that perform targeted updates through localized parameter modifications (Zhang et al., 2025; Scialanga et al., 2025; Li et al., 2025; Rozner et al., 2024). Existing knowledge editing methods roughly fall into two families: In-Place Parametric Edits (Meng et al., 2022a; Dai et al., 2025; Han et al., 2024; Fu et al., 2025; Jiang et al., 2025) that directly update model weights under the vanilla architecture, and External-Assisted Edits (Hartvigsen et al., 2023; Li et al., 2025; Tan et al., 2023; Wang et al., 2024d, a) that introduce auxiliary modules (e.g., hypernetworks or memory) to support updates at inference time. Despite encouraging results under single-edit and teacher-forced evaluations, their reliability in real-world autoregressive generation remains limited (Jiang et al., 2024; Tan et al., 2023; Chen et al., 2024; Gu et al., 2024a; Huang et al., 2024).
In this work, our empirical analysis reveals two recurring issues. (1) Pre-trained capability degradation. Many existing methods perform edits by maximizing the likelihood of the new target on a small set of editing samples. Without explicit constraints or regularization, this few-shot objective can induce over-specialization to the editing examples, degrading pre-trained capabilities. In lifelong editing, as illustrated in Figure 1 (left), such drift may accumulate across lifelong edits and may even lead to model collapse (Gupta et al., 2024, 2025b). (2) Parametric knowledge–behavior mismatch in autoregressive generation. More critically, Figure 1 (right) shows that even when the edited fact is stored parametrically, it may not be consistently expressed along autoregressive trajectories. A key reason is the train-test mismatch between teacher forcing and autoregressive inference: edits are optimized on ground-truth prefixes, whereas real-world generation conditions on self-generated prefixes, inducing a prefix distribution shift that makes edited behavior fragile and prone to reversion.
To address these limitations, we propose EtCon, a new edit-then-consolidate paradigm (Fig. 4) that couples constrained targeted edits with post-edit consolidation to improve both editing reliability and pre-trained capability preservation. Our training involves a two-stage pipeline: Stage 1 (Edit). Targeted Proximal Supervised Fine-Tuning (TPSFT), a refined variant of PSFT (Zhu et al., 2025), performs a constrained targeted edit by updating only knowledge-relevant FFN layers, with trust-region constraints and self-generated CoT-based targets. This design localizes parameter updates and mitigates over-specialization, thereby preserving pre-trained capabilities. Stage 2 (Consolidate). GRPO consolidates the edit via trajectory-level optimization under reward signals. This stage aligns the parametric update with inference-time behavior, improving the robustness of the edited knowledge along autoregressive trajectories. We conduct extensive experiments on three datasets with Llama-3-8B-Instruct and Qwen2.5-7B-Instruct. Under auto-regressive generation with natural stopping and an LLM-as-a-judge protocol (OpenAI, 2025), EtCon improves editing reliability and generalization by 35%-50% over strong baselines. It also significantly enhances locality while preserving critical pre-trained capabilities. Our contributions are as follows:
(1) We empirically show that incorporating a post-edit consolidation stage improves knowledge editing reliability under realistic scenarios.
(2) We propose EtCon, a novel knowledge editing paradigm: TPSFT for localized parametric edits, followed by GRPO for trajectory-level consolidation that aligns parametric knowledge with actual generation behavior.
(3) Extensive experiments demonstrate that EtCon improves editing reliability and generalization, while preserving pre-trained capabilities under realistic evaluation settings.
| Method | Reli. | Gener. | Local. | |
| Llama-3-8B-Instruct | Pre-Edit | 2.8 | 2.4 | 38.6 |
| Pre-Edit(+GRPO) | 5.2 (+2.4) | 4.7 (+2.3) | 38.4 (-0.2) | |
| FT-M | 16.6 | 15.5 | 29.3 | |
| FT-M(+GRPO) | 62.9 (+46.3) | 52.7 (+37.2) | 24.9 (-4.4) | |
| ALPHAEDIT | 18.7 | 14.0 | 6.3 | |
| ALPHAEDIT(+GRPO) | 50.4 (+31.7) | 38.7 (+24.7) | 5.4 (-0.9) |
2 Related Work
2.1 Overview of Knowledge editing methods
From the perspective of model architecture, the three paradigms introduced above can also be coarsely grouped into two families. In-Place Parametric Edits methods preserve the vanilla LLM architecture. The locate-then-edit paradigm (Meng et al., 2022a; Dai et al., 2025; Li et al., 2024; Zhong et al., 2025; Zhang et al., 2024c) identifies knowledge locations within LLMs and modifies targeted parameters through gradient-based or analytical solutions. PEFT methods (Zhu et al., 2020; Han et al., 2024; Wang et al., 2024c; Gupta et al., 2025a) directly update model parameters via regularized gradient descent to achieve knowledge updates while constraining side effects (Liu et al., 2025a). These approaches seamlessly integrate with existing deployment infrastructure without additional inference latency. External-Assisted Edits methods rely on auxiliary modules for knowledge modification. Meta-learning approaches (Tan et al., 2023; Hartvigsen et al., 2023; Li et al., 2025) train hypernetworks to generate parameter updates, while memory-based methods (Hartvigsen et al., 2023; Zhang et al., 2024b; Chen et al., 2024) encode knowledge in external modules that the LLM retrieves during inference. Despite their superior performance in balancing reliability and locality, external methods introduce deployment complexity. Given these trade-offs, our work advances In-Place Parametric Edits with a focus on more realistic lifelong knowledge editing scenarios.
2.2 Evaluation of Knowledge editing methods
Existing work (Fang et al., 2024; Qi et al., 2025; Scialanga et al., 2025) typically evaluates knowledge editing with three standard metrics. Reliability measures edit success as the fraction of cases where . Generalization applies the same criterion to rephrased queries about the edited knowledge. Locality assesses whether the edit preserves responses to related, unedited facts. In these protocols, inputs are often templated queries with minimal context, outputs are constrained to a fixed answer format (e.g., via truncation or in-context exemplars), and decoding commonly uses teacher forcing with ground-truth prefixes. Recent studies, however, have highlighted the fragility of such evaluation paradigms. Consequently, this paper adopts a realistic evaluation approach for knowledge editing methods. The full real-world evaluation framework is described in Appendix A.6.
3 Preliminary
3.1 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) (Guo et al., 2025) is a trajectory-level reinforcement learning algorithm that optimizes language model policies through relative reward comparisons within sampled groups. Given a policy and a dataset , GRPO maximizes expected reward while constraining deviation from a reference policy :
where is a reward function and controls regularization strength. The optimization employs a clipped surrogate objective:
where is the importance ratio, and is the group-relative advantage computed from trajectories sampled for the same query . This design eliminates the need for a separate critic network while enabling stable policy updates. Notably, GRPO’s trajectory-level optimization makes it naturally suited for aligning knowledge that is parametrically encoded but not reliably applied during autoregressive generation.
3.2 Motivation for Knowledge Consolidation
Our key insight is that successful knowledge editing requires a two-stage process rather than parametric knowledge editing alone. Specifically, recent studies reveal a critical gap: while existing methods achieve high success under teacher-forcing evaluation, they can fail in realistic autoregressive generation. Through systematic investigation, we posit that an explicit knowledge consolidation stage is crucial for mitigating this gap. Without a knowledge consolidation stage, edited knowledge may remains superficially encoded at the parametric level, failing to propagate to the model’s generation behavior. To validate this hypothesis, we augment existing editing methods with GRPO as a post-editing consolidation step. As shown in Table 3 and Figure 3, introducing consolidation dramatically improves both reliability and generalization across multiple baselines. Furthermore, applying GRPO directly to unedited models yields minimal gains, confirming that consolidation requires prior parametric editing as a foundation. These findings establish that parametric knowledge updates and behavioral alignment are complementary but distinct requirements for successful knowledge editing.
4 EtCon
In this section, we present the EtCon framework (as shown in Figure 4). Stage I employs Targeted Proximal Supervised Fine-Tuning (TPSFT) to perform localized knowledge editing with PPO-style ratio clipping that constrains policy divergence, thereby limiting spillover while preserving pre-trained abilities (§4.1); Stage II applies GRPO with a comprehensive reward to consolidate edited knowledge at the trajectory level under autoregressive generation (§4.2).
4.1 Knowledge Editing via Targeted Proximal Fine-Tuning
In this section, we introduce Targeted Proximal Supervised Fine-Tuning (TPSFT) as a refined knowledge-editing method designed to improve editing efficacy while better preserving pre-trained capabilities. This approach differs from raw PSFT that updates the whole LLMs, we restrict updates exclusively to specific FFN layers, widely recognized as the primary stores of factual knowledge (Meng et al., 2022a). This targeted update strategy effectively injects new knowledge without modifying the model architecture, while minimizing disruption to pre-trained capabilities.
We consider a knowledge editing instance , where is the input query and is the target answer. Let parameterized by denote the vanilla LLM before the current edit. We partition the current parameters into two disjoint sets: the trainable FFN parameters and the frozen parameters , such that . The objective of TPSFT is to optimize to learn the updated parameters , yielding the post-edit policy where .
A distinctive aspect of TPSFT is the use of self-generated CoT-augmented training labels. For each instance , we: (1) prompt the vanilla LLM to generate a CoT reasoning path for (see Appendix A.5), and (2) replace the generated answer with , yielding the training target . This design integrates the CoT into the supervision signal. Rather than overfitting to a sharp, isolated input-output mapping, the model learns to associate the new knowledge with valid reasoning patterns. To ensure data quality, we filter out generated chains that are logically incompatible with the target (see Appendix A.5), thereby preserving the model’s intrinsic generation capabilities while injecting the new fact.
Formally, given the editing dataset , for each token position in the target sequence , we define the state and the target token . We optimize by minimizing the following TPSFT loss:
Here, is a hyperparameter that defines the clipping threshold. Following Zhu et al. (2025), the ratio clipping mechanism provides implicit trust-region regularization by (i) bounding the log-probability change per sample to , and (ii) nullifying gradients when the ratio exceeds the upper bound. The probability ratio is the core of this objective and is defined as:
where is the probability of generating the target token given the context from the model with the current FFN parameters, and is the corresponding probability from the reference policy. At the start of the editing process, this reference policy is the initial vanilla LLM. For each subsequent edit instance in the sequential editing process, it is then updated to be the state of the model resulting from the immediately preceding edit.
This objective function acts as a trust-region regularizer via ratio clipping. The term encourages the model to increase the likelihood of the target answer, similar to standard supervised fine-tuning. Crucially, when this ratio exceeds the threshold , the gradient with respect to is nullified. This design effectively removes optimization pressure from well-learned samples, constraining parametric deviation and preventing individual edits from dominating the parameter updates.
However, despite these constraints, TPSFT remains a teacher-forced training mechanism implicitly conditioned on ground-truth prefixes . During inference, the model must rely on its own generated history . This discrepancy creates a fundamental train-inference mismatch: while the knowledge is parametrically encoded, it may not be reliably elicited during autoregressive generation. This critical limitation necessitates the subsequent consolidation phase.
4.2 Knowledge Consolidation via GRPO
As motivated in § 3.2, parametric editing alone is insufficient—edited knowledge must be consolidated at the behavioral level. Building on GRPO’s trajectory-level optimization (§ 3.1), we design a consolidation algorithm that transforms parametrically encoded knowledge into reliably generated outputs.
Starting from the edited model obtained via TPSFT, we optimize the policy (initialized as ) through autoregressive sampling on the editing dataset . For each query , the model generates complete trajectories , which are then scored against the target answer . The group-relative advantage formulation enables the model to learn from comparative trajectory quality without requiring absolute reward calibration. The reference policy is set to , ensuring that consolidation refines generation behavior while preserving the parametric edits from Stage I.
| ZsRE | COUNTERFACT | QAEdit | |||||||||||
| Method | Reli. | Gen. | Loc. | Avg. | Reli. | Gen. | Loc. | Avg. | Reli. | Gen. | Loc. | Avg. | |
| Qwen2.5-7B -Instruct | Pre-edit | 4.4 | 3.2 | 28.5 | 12.0 | 1.0 | 0.5 | 36.9 | 12.8 | 9.8 | 10.1 | 36.2 | 18.7 |
| FT-M | 5.6 | 5.5 | 23.1 | 11.4 | 3.2 | 3.1 | 24.4 | 10.2 | 14.6 | 14.5 | 30.7 | 19.9 | |
| MEMIT | 0.0 | 0.1 | 0.0 | 0.0 | 0.0 | 0.2 | 0.1 | 0.1 | 0.4 | 0.3 | 0.2 | 0.3 | |
| ALPHAEDIT | 15.9 | 11.5 | 6.8 | 11.4 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
| WISE | 4.5 | 3.3 | 19.1 | 9.0 | 1.4 | 1.5 | 31.0 | 11.3 | 7.1 | 9.7 | 16.9 | 11.2 | |
| GRACE | 4.8 | 4.7 | 28.1 | 12.5 | 1.9 | 1.2 | 36.8 | 13.3 | 14.2 | 13.0 | 36.0 | 21.1 | |
[…]
| Llama-3 -8B-Instruct | Pre-edit | 2.8 | 2.4 | 38.6 | 14.6 | 0.6 | 0.8 | 31.8 | 11.1 | 12.7 | 12.5 | 44.3 | 23.2 |
| FT-M | 16.6 | 15.5 | 29.3 | 20.5 | 27.9 | 18.6 | 10.5 | 19.0 | 34.1 | 33.2 | 30.1 | 32.5 | |
| MEMIT | 0.1 | 0.1 | 0.0 | 0.1 | 0.3 | 0.7 | 0.4 | 0.5 | 0.2 | 0.7 | 0.0 | 0.3 | |
| ALPHAEDIT | 18.7 | 14.0 | 6.3 | 13.0 | 61.0 | 43.8 | 16.1 | 40.3 | 18.2 | 14.9 | 7.5 | 13.5 | |
| WISE | 4.3 | 3.1 | 2.2 | 3.2 | 1.3 | 0.8 | 31.3 | 11.1 | 8.1 | 13.3 | 0.9 | 7.4 | |
| GRACE | 5.5 | 3.9 | 37.9 | 15.8 | 2.1 | 0.7 | 32.1 | 11.6 | 15.4 | 13.4 | 34.2 | 21.0 | |
[…]
The reward function evaluates multiple aspects of the generated trajectory (as illustrated in Figure 4):
where measures factual accuracy (whether the final answer matches ), enforces task-specific output format requirements, encourages concise outputs without extraneous tokens, and rewards internal reasoning coherence and alignment between intermediate steps and the final answer.
This consolidation step integrates the parametric knowledge acquired through TPSFT into the model’s inference policy. By optimizing trajectory-level generation under autoregressive sampling rather than teacher-forcing, GRPO mitigates the train-inference mismatch, ensuring that edited knowledge is reliably elicited in actual behavior.
5 Experiments
5.1 Experiment Settings
Datasets and Models. This work utilizes 1000 samples from each of four benchmark, ZsRE (Levy et al., 2017), COUNTERFACT (Meng et al., 2022a), MQuAKE-CF-v2 (Zhong et al., 2023) (1-edit subset), and QAEdit (Yang et al., 2025b), to evaluate the performance on knowledge editing tasks. We focus on the lifelong editing with sequential edits (batch size 1), which better reflects realistic deployment scenarios where knowledge updates arrive continuously over time. We select the Llama-3-8B-Instruct (Dubey et al., 2024) and Qwen2.5-7B-Instruct (Qwen et al., 2025), as the base models for editing. For general ability evaluation, we use C-Eval (Huang et al., 2023), CoQA (Reddy et al., 2019), DROP (Dua et al., 2019), SQuAD 2.0 (Rajpurkar et al., 2018) and LogiQA (Liu et al., 2020).
Baselines. We compare our method against two main categories: In-Place Parametric Edits methods (FT-M (Zhang et al., 2024a), MEMIT (Meng et al., 2022b), ALPHAEDIT (Fang et al., 2024), MMKE (Fu et al., 2025)) and External-Assisted Edits methods (WISE (Wang et al., 2024a), GRACE (Hartvigsen et al., 2023)). In-Place Parametric Edits methods are the main focus of this work, and we select the most representative methods as baselines.
Implementation Details. We conduct experiments using EasyEdit (Xu et al., 2025) for evaluating various baselines, and employ the lm-evaluation-harness (Gao et al., 2024a) for assessing general capabilities. TPSFT is implemented through PSFT (Zhu et al., 2025) for edit stage, while GRPO is built upon the EasyR1 (Zheng et al., 2025b) for the consolidation stage. More details are in Appendix A.1.
Evaluation Metrics. We evaluate our method along two principal axes: editing performance and pre-trained capability preservation. To assess editing performance, we employ the LLM-as-judge framework from (Yang et al., 2025b; Gao et al., 2024b; Gu et al., 2024b; Wang et al., 2024b), which mitigates the overestimation issue inherent in token-based metrics. Note that while GRPO training uses efficient rule-based rewards (§4.2), evaluation employs GPT-4.1 for semantic assessment to ensure rigor. Specifically, we conduct a binary (correct/incorrect) evaluation of the model’s edited outputs. For ZsRE, COUNTERFACT, and QAEdit, we measure Reliability (edit success), Generalization (effectiveness on related inputs), and Locality (impact on unrelated inputs). For MQuAKE-CF-v2, we report Edit-wise accuracy and Multi-hop accuracy. We report pretrained capability on C-Eval and LogiQA, and Exact Match (EM) and F1 on CoQA, DROP, and SQuAD 2.0. Real-world evaluation details are provided in Appendix A.6.
5.2 Main Results
Edit Efficiency. EtCon consistently achieves the best overall performance under real-world lifelong editing evaluation across benchmarks and base models. Table 1 presents comprehensive results on two LLMs. Across all six settings (3 benchmarks 2 base models), EtCon attains the top Avg. score, with large margins over the strongest competing method. On Qwen2.5-7B-Instruct, EtCon reaches Avg. of 51.5/44.2/56.8 on ZsRE/COUNTERFACT/QAEdit (improving the runner-up by +39.0/+30.9/+35.7). On Llama-3-8B-Instruct, EtCon further achieves Avg. of 55.6/48.2/55.7 (+35.1/+7.9/+23.2). These gains are primarily driven by substantially higher Reliability and Generalization, while maintaining competitive Locality (24.2%–33.6%), indicating that EtCon can internalize new facts and generalize them to paraphrases under sequential edits without overly compromising unrelated behaviors. We attribute this robustness to our two-stage design: TPSFT performs localized, stability-aware updates, and GRPO subsequently consolidates the edited knowledge under autoregressive trajectories.
| Editing Performance | General Capabilities | ||||||||||||
| Method | Reli. | Gen. | Loc. | Avg. | C-Eval | CoQA | DROP | SQuAD | LogiQA | Avg. | |||
| Acc. | EM | F1 | EM | F1 | EM | F1 | Acc. | ||||||
| Pre | 9.8 | 10.1 | 36.2 | 18.7 | 79.5 | 54.5 | 70.1 | 2.2 | 9.9 | 9.9 | 18.9 | 38.7 | 38.0 |
| FT-M | 14.6 | 14.5 | 30.7 | 19.9 | 75.9 | 21.3 | 38.7 | 2.4 | 13.3 | 2.9 | 11.2 | 37.0 | 25.3 |
| FT-M+Con. | 42.3 | 34.1 | 31.9 | 36.1 | 77.0 | 26.2 | 46.6 | 2.8 | 14.6 | 4.4 | 13.5 | 37.8 | 27.9 |
| MMKE | 12.2 | 10.4 | 34.2 | 18.9 | 79.3 | 60.3 | 74.6 | 10.3 | 24.3 | 13.8 | 21.1 | 41.0 | 40.6 |
| MMKE+Con. | 37.2 | 31.4 | 31.0 | 33.2 | 78.8 | 59.1 | 73.3 | 8.5 | 21.6 | 12.1 | 19.6 | 39.2 | 39.0 |
| ALPHA | 0.0 | 0.0 | 0.0 | 0.0 | 23.0 | 0.0 | 0.0 | 0.0 | 0.0 | 50.1 | 50.1 | 21.8 | 18.1 |
| ALPHA+Con. | 0.0 | 0.0 | 0.0 | 0.0 | 23.0 | 0.0 | 0.0 | 0.0 | 0.0 | 50.1 | 50.1 | 21.8 | 18.1 |
[…]
| Methods | Edit-wise | 2-hop | 3-hop | 4-hop |
| Base | - | 69.1 | 36.2 | 85.4 |
| FT-M | 19.1 | 2.7 | 3.8 | 4.2 |
| ALPHA | 20.4 | 13.9 | 11.3 | 9.5 |
| MMKE | 9.3 | 12.0 | 7.1 | 22.9 |
| MEMIT | 0.4 | 1.1 | 1.3 | 0.0 |
| EtCon | 54.1 | 24.2 | 14.6 | 33.3 |
| Stage | Methods | Reli. | Gen. | Loc. | GC-Avg |
| Base | - | 0.6 | 0.8 | 31.8 | 52.85 |
| Edit. | w/ SFT | 1.4 | 0.3 | 30.7 | 50.31 |
| w/ TPSFT | 3.3 | 1.8 | 30.2 | 54.40 | |
| Con. (after TPSFT) | w/o | 56.1 | 22.4 | 24.7 | 50.17 |
| w/o | 51.6 | 27.2 | 25.1 | 49.75 | |
| Complete | 67.1 | 53.4 | 24.2 | 50.25 |
| Edit Scope | Reli. | Gen. | Loc. | GC-Avg |
| Full Param. | 47.3 | 41.5 | 35.8 | 58.17 |
| Target edit | 75.1 | 63.0 | 32.3 | 55.83 |
Scaling to Longer Editing. Figure 5 shows performance evolution when scaling to 3,000 sequential edits on ZsRE. EtCon maintains stable performance: Reliability and Generalization decrease moderately from 73.5% to 58.2% and 63.1% to 48.7% respectively, while Locality remains within a narrow band (28%–32%). In contrast, FT-M degrades rapidly, with Reliability dropping from 16.6% to near-zero. These results demonstrate that EtCon remains stable even under longer editing sequences.
6 Analysis
6.1 Effectiveness of Consolidation Stage
Overall Performance of EtCon. As shown in Table 2, EtCon achieves the best trade-off between editing effectiveness and general capability preservation. EtCon attains 75.1% Reliability and 63.0% Generalization, substantially outperforming all baselines. Meanwhile, EtCon maintains general capabilities close to the original model (C-Eval: 78.5 vs 79.5, CoQA F1: 69.4 vs 70.1), demonstrating that our two-stage design successfully balances knowledge injection with pre-trained capability preservation.
Performance on Multi-hop Task. Table 4 further demonstrates EtCon’s advantage on MQuAKE-CF-v2. EtCon achieves 24.2% and 33.3% accuracy on 2-hop and 4-hop queries respectively, substantially outperforming FT-M (2.7%/4.2%). This indicates that our consolidation stage can propagate localized edits into more consistent multi-step inference. Under the EtCon paradigm, incorporating explicit reward for multi-hop queries during consolidation could further strengthen performance on multi-hop tasks.
Universality of Knowledge Consolidation. To evaluate the generality of our knowledge consolidation stage, we augment three baselines (FT-M, MMKE, and ALPHAEDIT) with Consolidation and compare them with EtCon on QAEdit. As shown in Table 2, adding Consolidation to FT-M and MMKE improves Reliability by 25% to 28% and Generalization by about 20%. These results suggest that Consolidation bridges the gap between edited parametric knowledge and autoregressive generation behavior, enabling more reliable knowledge utilization in practice.
Preservation of Pre-trained Capabilities. Evaluations across benchmarks show that the knowledge consolidation stage is non-destructive and preserves general capabilities. However, it cannot recover capabilities lost during editing: once ALPHAEDIT collapses, the knowledge consolidation stage cannot restore it. This indicates that reliable lifelong editing requires both lossless knowledge injection and post-edit consolidation. Future work under the EtCon paradigm should prioritize lossless knowledge editing.
6.2 Ablation Study
Edit Stage Ablation. We compare TPSFT with standard SFT in Table 5. The results show that neither method alone yields reliably editable behavior, as evidenced by consistently low Reliability and Generalization. Nevertheless, TPSFT offers a clear advantage in capability preservation, substantially reducing the degradation observed with SFT.
Consolidation Stage Ablation. We ablate key components of our consolidation reward (Table 5). Removing causes a clear drop (Reliability , Generalization ); inspection shows this encourages “reward hacking,” where the model emits extraneous content (e.g., both old and new facts) to inflate scores. Removing is even more detrimental to Reliability (), often producing catastrophic self-contradictions: the model states the correct answer and then immediately reverts. These ablations suggest that both reward components help mitigate reward hacking and improve output coherence; see Appendix A.8 for case studies.
Parameter Sensitivity. Figure 4 analyzes TPSFT hyper-parameters. Clipping threshold : Increasing improves Reliability and Generalization but reduces Locality, with a sharp drop at . Thus, provides the best trade-off, as larger allows stronger updates but increases policy drift. Reference update interval: Larger intervals consistently reduce Reliability and Generalization, while Locality remains nearly unchanged. Frequent updates (Interval) perform best because stale references increase the policy gap and over-activate clipping, limiting effective updates.
Effect of CoT Augmentation. Figure 6 compares TPSFT with and without CoT-augmented training labels. Without CoT augmentation, Reliability is slightly higher, but Locality drops sharply and general capabilities degrade (e.g., lower CoQA compared to base). This is because supervising only on the few target tokens leads to overfitting. CoT augmentation encourages the model to learn valid reasoning patterns rather than overfitting to isolated input-output mappings, thereby better preserving general capabilities.
Efficiency of Target Edit. We compare Target Edit, which updates a small set of targeted FFN weights, with Full Param update under the same TPSFT setup. As shown in Table 6, Target Edit yields substantially higher Reliability and Generalization, while incurring only a slight drop in general capability (GC-Avg). In contrast, Full Param attains notably weaker editing performance under the same optimization budget. We attribute this gap to reduced optimization efficiency under ratio clipping. With limited update steps per edit, Full Param spreads gradients across the whole model, weakening the learning signal on the edited fact. In addition, clipping may saturate early for many tokens, further shrinking effective updates. Overall, restricting TPSFT provides a more controllable and effective mechanism for knowledge editing.
7 Conclusion
In this work, we propose EtCon, an edit-then-consolidate framework for reliable knowledge editing. In Edit Stage, Targeted Proximal Supervised Fine-Tuning (TPSFT), performs constrained edits to inject knowledge while preserving pre-trained capabilities. In the Consolidate Stage, Group Relative Policy Optimization (GRPO), consolidates edits by aligning autoregressive generation with updated facts. Extensive experiments demonstrate that EtCon achieves superior editing reliability and generalization while preserving pre-trained capabilities in real-world editing scenarios.
Appendix A Appendix
A.1 Additional Implementation Details
Baseline Configuration. For our baseline experiments, we utilize the EasyEdit framework. All hyperparameters adhere to the default configurations of the respective comparison methods, with further details provided in (Yang et al., 2025b; Qi et al., 2025).
Editing Stage. For our proposed EtCon method, we update only the FFN down-projection layers (mlp.down_proj) in layers 7–11 of Llama-3-8B-Instruct and layers 5–9 of Qwen2.5-7B-Instruct, following prior works (Zhang et al., 2024a; Meng et al., 2022a; Geva et al., 2021; Qi et al., 2024) on knowledge editing. Following the lifelong editing paradigm, we use AdamW with learning rate and set of TPSFT, sequentially performing edits (batch size 1) on the same model. TPSFT is trained for 5 epochs with at most 6 update steps per edit using early stopping.
Consolidation Stage. We optimize the inference-time policy with GRPO. The comprehensive reward function in Equation (5) uses the following weight coefficients: for , for , for , and for . These weights were determined through extensive empirical experiments to balance factual accuracy with output quality. All reward components are binary (0/1), with computed via exact string matching between the extracted answer and the ground-truth after normalization. All specific hyperparameters are available in Table 7.
| Configuration | Value |
| Model Configuration | |
| Precision | BFloat16 |
| Max Prompt Length | 2k |
| Max Response Length | 2k |
| Training Hyperparameters | |
| Learning Rate | |
| Optimizer | AdamW (BF16 variant) |
| Global Batch Size | 64 |
| Rollout Batch Size | 256 |
| Micro Batch Size (Update) | 4 |
| Micro Batch Size (Experience) | 16 |
| Training Step | 100 |
| Gradient Clipping | 1.0 |
| Rollout Configuration | |
| Number of Rollouts () | 8 |
| Temperature | 1.0 |
| Top-p | 0.99 |
| Infrastructure | |
| GPUs | 8 × NVIDIA H800 |
| Tensor Parallelism | 1 |
| FSDP | Enabled |
| CPU Offloading | Disabled |
| Gradient Checkpointing | Enabled |
| Validation | |
| Validation Batch Size | 512 |
| Validation Frequency | Every 5 steps |
| Validation before Training | Yes |
A.2 Additional Results
A.2.1 Effect of Edited Layers on EtCon
| Layers | Reli. | Gener. | Local. |
| 7-8-9-10-11 | 73.5 | 63.1 | 30.2 |
| 12-13-14-15-16 | 78.1 | 58.3 | 24.1 |
| 17-18-19-20-21 | 76.7 | 53.2 | 17.3 |
Experimental Setup. We further investigate the impact of editing different FFN layers on EtCon’s performance through an ablation study, as shown in Table 8 and Figure 7.
Main Finding. We observe that under identical hyperparameter settings, editing early layers (Layers 7-11) outperforms deeper layers (Layers 17-21) in both locality and generalization.
Reward Hacking Phenomenon. Analyzing the “high reward, low performance” phenomenon in deeper layers shown in Figure 7, we find that editing deeper layers is more prone to triggering Reward Hacking.
Mechanistic Interpretation. Existing mechanistic interpretability research suggests that shallow layers mainly store factual knowledge while deeper layers handle information integration and reasoning. Based on this, we posit that the observed phenomenon stems from a misalignment between retained knowledge and updated information.
Cognitive Conflict Analysis. Editing only deeper layers may cause retained knowledge in shallow layers to conflict with the injected knowledge in deeper layers. Faced with this cognitive conflict, the LLM likely adopts a speculative strategy to maximize rewards. This reward hacking leads to internal model confusion, thereby causing a decline in performance.
A.2.2 Portability Evaluation
Superior Cross-Architecture Portability. As shown in the Figure 8, EtCon demonstrates superior portability across architectures, achieving 26.2% on Llama-3 and 24.3% on Qwen2.5, consistently outperforming the strongest baseline FT-M (18.8% and 21.8%).
Consolidation Effectiveness. This validates that the consolidation stage effectively enables the model to utilize edited facts for reasoning.
Limitations and Future Directions. However, the absolute gains remain limited compared to single-hop reliability. We attribute this to EtCon being trained solely on single-hop QA without explicit supervision for multi-hop compositionality; thus, incorporating such supervision is a natural next step to further enhance portability.
A.3 Evaluation on Reasoning-Oriented Architectures
| Layers | Reli. | Gener. | Local. |
| 5-6-7-8-9 | 88.6 | 53.5 | 17.0 |
| 13-14-15-16-17 | 79.6 | 42.6 | 16.1 |
| 23-24-25-26-27 | 83.0 | 52.9 | 7.3 |
Experimental Setup. To assess the robustness of our method on architectures with inherent reasoning capabilities, we extended our evaluation to DeepSeek-R1-Distill-Qwen-7B using 1,000 samples from the ZsRE. Following our main experiments, we target the MLP down-projection weights (mlp.down_proj) for knowledge editing.
Shallow Layers Yield Best Trade-off. As reported in Table 9, EtCon maintains high editing efficacy on this reasoning-oriented model. In particular, editing the shallow layers (Layers 5–9) yields the best trade-off, achieving 88.6% Reliability and 53.5% Generalization while preserving acceptable Locality (17.0%). This suggests that our Edit-then-Consolidate paradigm is compatible with the model’s intrinsic reasoning processes rather than disrupting them.
Deep-Layer Shortcut Phenomenon. The green curve exhibits the fastest initial convergence (Steps 0-20). Since these layers are proximal to the output projection, the model can quickly maximize the reward by establishing a direct mapping to the target answer. However, similar to the phenomenon observed in Appendix A.2.1, this configuration plateaus at a lower reliability level. This suggests that modifying only the deep layers could lead the model to adopt a ”shortcut” strategy which may leave the model vulnerable to the cognitive conflict between retained shallow knowledge and updated deep injections.
Shallow-Layer Constructive Consolidation. In contrast, the red curve displays a distinct ”warm-up” phase (Steps 0-15) followed by a sustained ascent to the highest performance plateau. This trajectory indicates that injecting knowledge into the shallow layers requires more optimization steps for the consolidate stage to propagate the changes through autoregressive generation pipeline. Crucially, this delay reflects a constructive consolidation process: the model is realigning its actual generation behavior with the updated parametric knowledge base. This results in convergence stability and reliability, demonstrating that our Edit-then-Consolidate paradigm is compatible with the intrinsic reasoning mechanisms of complex architectures.
A.4 Time Efficiency Analysis
| Methods | TPSFT (Ours) | AlphaEdit | MEMIT | GRACE | WISE | FT-M |
| Avg. Time / Edit | 6.01s | 7.39s | 7.78s | 3.02s | 2.68s | 0.61s |
In our earlier analysis, we argued that reliable knowledge editing in realistic settings requires both a knowledge editing stage and a knowledge-consolidation stage, and that any speed advantage should be predicated on reliability. Accordingly, in our runtime study, we report the computational cost for these two stages separately.
Editing stage. We first evaluate the average editing latency of different methods on Qwen2.5-7B-Instruct using the QAE dataset. For each method, we perform a sequence of 100 single-instance edits and measure the wall-clock time required to complete each individual edit. All methods are run under the same configuration. The results are summarized in Table 10. TPSFT attains an average editing time of 6.01 seconds, which is comparable to that of AlphaEdit (7.39 s) and MEMIT (7.78 s). GRACE (3.02 s) and WISE (2.68 s) exhibit lower latency, while FT-M yields the lowest runtime (0.61 s). Overall, TPSFT incurs a per-edit cost that is on par with representative parameter-editing baselines, indicating that its computational overhead at this stage remains moderate and compatible with practical deployment.
Consolidation stage. Starting from edited LLM by FT-M, AlphaEdit, MMKE, and TPSFT, we apply GRPO as the consolidation algorithm on QAE using the same hyperparameter configuration for all methods and train for 15 steps (corresponding to approximately one hour of wall-clock time). Figure 10 plots the comprehensive reward curves. TPSFT+Con (our method) exhibits steadily increasing rewards and is close to convergence by step 15, whereas FT-M+Con and MMKE+Con also improve during training but converge noticeably more slowly; in contrast, the curve of AlphaEdit+Con remains essentially flat, as the model has already collapsed in the editing stage. These results indicate that, under the same configuration, TPSFT+Con attains faster convergence in the consolidation stage than the baselines.
A.5 CoT Generation and Processing
To generate Chain-of-Thought reasoning paths during TPSFT training, we employ the following prompt template that encourages natural reasoning while ensuring convergence to target answers:
Design Rationale: This prompt template serves three critical functions in our TPSFT implementation:
1.
Natural Reasoning Preservation: By requesting step-by-step analysis without explicitly mentioning knowledge updates, the model generates reasoning paths consistent with its pre-trained style.
2.
Target Alignment: Providing the target answer guides the generation toward correct conclusions while allowing flexibility in reasoning approaches.
3.
Structured Output: The answer tag format ensures clean extraction and replacement during training data preparation, while the reasoning portion provides the smooth distribution over trajectories discussed in Section 4.1.
After generation, we enclose the reasoning path within <think></think> tags to explicitly demarcate the thought process, and overwrite the content inside the <answer></answer> tags with the verified new target fact, yielding training labels that combine natural reasoning patterns with an exact gold answer. We additionally discard CoT samples whose final answer is clearly inconsistent with the ground-truth and regenerate them, further reducing the risk of noisy supervision.
A.6 Real-world evaluation Details
In this work, we follow the design (Yang et al., 2025b) and use the better reflects real-world application scenarios evaluation to comprehensively measure the performance of knowledge editing methods. Specifically, our evaluation process consists of three key stages:
(1) For Input: To assess the model’s ability to deeply integrate and apply new knowledge, our inputs include both factual questions and instructions that require multi-step reasoning. This challenges the model to go beyond mechanically recalling the edited information and instead perform logical deductions based on it. For this purpose, we use the system prompt: Please reason step by step, then answer {question}.
(2) For Output: For the edited model output, we use the model’s complete auto-regressive generation as the object of evaluation, up to its predefined stop token. This approach allows us to assess not only the accuracy of the answer but also to examine the post-edit model’s performance in aspects such as fluency, coherence, and whether it introduces irrelevant content.
(3) LLM-as-a-Judge Evaluation: To achieve a scalable and objective evaluation, we introduce a more powerful Large Language Model (LLM) to act as a ”judge.” This judge model makes its decision by comprehensively considering the original question, the ground-truth answer (Target), and the full answer content from the edited model, ultimately providing a binary (correct/incorrect) judgment. The full judge prompts are provided in Figure 11 and Figure 12
A.7 Metrics for MQuAKE-CF-v2
For multi-hop task evaluation, we sample 1000 instances from MQuAKE-CF-v2 proportionally across 2-hop, 3-hop, and 4-hop questions. We report the following metrics:
Edit-wise accuracy: measures whether the model correctly answers direct queries about edited facts.
Multi-hop accuracy (2/3/4-hop): evaluates whether the model can perform multi-step reasoning that involves the edited knowledge.
A.8 Analysis of Reward Hacking Cases
Analysis of Reward Hacking Patterns: The two cases in Figures. 13 and 14 reveal distinct failure modes in the absence of proper reward design. In Figure 13, the model exhibits ”self-correction” behavior—correctly reasoning through the problem but then artificially inserting the target answer followed by an immediate correction. This pattern emerges when Rconsistency is absent, as the model attempts to maximize accuracy rewards without maintaining logical coherence. Figure 14 demonstrates ”answer hedging” where the model provides multiple answers to maximize the probability of including the correct one. This occurs without Rcleanliness, as there’s no penalty for extraneous content. These cases underscore that comprehensive reward design is not merely beneficial but essential for preventing models from exploiting loopholes in the optimization objective. The 11.0% and 15.5% performance drops observed when removing these reward components (Table 5) quantitatively confirm their critical role in maintaining robust consolidation.